Circular No. 9099

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V496 SCUTI R. J. Rudy and T. R. Prater, The Aerospace Corporation; R. C. Puet-

ter, University of California at San Diego; and R. B. Perry, NASA, report 0.9- to 2.5-μm spectroscopy of V496 Sct (cf. IAUCs 9093, 9097) on Nov. 27.08 UT using the Aerospace Corporation’s Near-Infrared Imaging Spectrograph on the 3-m Shane reflector at Lick Observatory. The nova displays very strong emission from the first overtone of carbon monoxide; this is a rare and extremely short-lived feature in novae and occurs just before dust formation. The nova also displays several C I lines that are among the brightest emission lines in the spectrum. The strongest of these display P-Cyg profiles, behavior not observed previously in novae. Very strong C I emission in early novae spectra is a common occurrence in novae that form dust. The authors believe that dust formation in V496 Sct is almost certain and that there is a good probability that it will occur within the next week. K. Baker helped in acquiring the measurements.

S/2009 (317) 1 W. J. Merline, Southwest Research Institute (SwRI); P. M. Tam-

blyn, Binary Astronomy, Dillon, CO, U.S.A., and SwRI; J. D. Drummond, Starfire Optical Range, Air Force Research Laboratory; J. C. Christou, Gemini Observatory; A. R. Conrad, W. M. Keck Observatory; B. Carry, Observatoire de Paris; C. R. Chapman, SwRI; C. Dumas, European Southern Observatory (Chile); D. D. Durda, SwRI; W. M. Owen, Jet Propulsion Laboratory; and B. L. Enke, SwRI, report the discovery on Nov. 24.4 UT of a satellite of minor planet (317) Roxane from K_{p}-band imaging using the 8-m Gemini-North telescope (+ Altair/NIRI adaptive-optics system). On Nov. 24.39992, the satellite was at separation 0′′.27 (projected separation 245 km) and position angle 76^{o}. The satellite was imaged in K_{p}-, H-, and J-bands and was tracked for more than 26 hours. The brightness difference in the K_{p} band is ∼ 2.7 mag, giving an estimated diameter (primary diameter assumed of 19 km) of the satellite of 5 km. This binary has characteristics that are very similar to the other wide binaries previously reported by the authors (see IAUCs 7827, 8075, 8232, 8293, 8297), all being consistent with the EEB formation mechanism of Durda et al. (2004, Icarus 170, 243). Among binaries that can be resolved by imaging, this appears to be the first of type E. Given the recent detection of dual lightcurve periods in (1509) Esclangona (CBET 2020), it is possible that separate spin periods could be extracted from lightcurve data of (317).

2009 December 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green